ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 220, 2017

Aoyama, K. and M. Toyoda Approximation of zeros of accretive operators in a Banach space

803

Artigiani, M. Exceptional ergodic directions in Eaton lenses  29

Bate, D., M. Csörnyei and B. Wilson The Besicovitch–Federer projection theorem is false in every infinite-dimensional Banach space

175

Buss, A., R. Exel and R. Meyer Reduced C*-algebras of Fell bundles over inverse semigroups

225

Calviño-Louzao, E., M. Fernández-López, E. García-Río and R. Vázquez-Lorenzo Homogeneous Ricci almost solitons

531

Cameron, P. J., M. Gadouleau, J. D. Mitchell and Y. Peresse Chains of subsemigroups

479

Chen, J., P. Li, S. K. Sahoo and X. Wang On the Lipschitz continuity of certain quasiregular mappings between smooth Jordan domains

453

Chen, W. Homology of braid groups, the Burau representation, and points on superelliptic curves over finite fields

739

Csörnyei, M. See D. Bate

De Bernardi, C. A. Extreme points in polyhedral Banach spaces  547

Exel, R. See A. Buss Farb, B. and S. Hensel Moving homology classes in finite covers of

graphs

605

Ferber, A., G. Kronenberg and E. Long Packing, counting and covering Hamilton cycles in random directed graphs

57

Fernández-López, M. See E. Calviño-Louzao Ferreira, L. C. F. and J. E. Pérez-López On the theory of Besov–

Herz spaces and Euler equations

283

Frenkel, P. E. Polynomial identities for matrices over the Grassmann algebra

791

Gadouleau, M. See P. J. Cameron Gao, N., V. G. Troitsky and F. Xanthos Uo-convergence and its

applications to Cesàro means in Banach lattices

649

García-Río, E. See E. Calviño-Louzao Garti, S., Y. Hayut and S. Shelah On the verge of inconsistency:

Magidor cardinals and Magidor filters 89 Gordienko, A., G. Janssens and E. Jespers Semigroup graded

algebras and graded PI-exponent

387

Greenfeld, B. Prime and primitive algebras with prescribed growth types

161

Hayut, Y. See S. Garti Hensel, S. See B. Farb Janssens, G. See A. Gordienko Jespers, E. See A. Gordienko Kaya, B. The complexity of the topological conjugacy problem for

Toeplitz subshifts

873

Konieczny, J. Combinatorial properties of Nil–Bohr sets  333

Kronenberg, G. See A. Ferber

Lambie-Hanson, C. Covering properties and square principles  617

Li, P. See J. Chen Long, E. See A. Ferber Malliaris, M. and S. Shelah Model-theoretic applications of cofi-

nality spectrum problems

947

M^{c}Carthy, J. E. and O. M. Shalit Spaces of Dirichlet series with the complete Pick property

509

Meyer, R. See A. Buss Mitchell, J. D. See P. J. Cameron Müller, C. and L. Speigelhofer Normality of the Thue–Morse

sequence along Piatetski-Shapiro sequences, II

691

Ngô, Q. A. and V. H. Nguyen Sharp reversed Hardy–Littlewood– Sobolev inequality on R^{n}

189

Nguyen, V. H. See Q. A. Ngô Ojala, T., V. Suomala and M. Wu Random cutout sets with spa-

tially inhomogeneous intensities

899

Pérez-López, J. E. See L. C. F. Ferreira Peresse, Y. See P. J. Cameron Procházka, A. and L. Sánchez-González Low distortion embed-

dings of some metric graphs into Banach spaces

927

Quoirin, H. R. and K. Umezu An indefinite concave-convex equation under a Neumann boundary condition I

103

Rivoal, T. and J. Roques Holomorphic solutions of E-operators  275

Roques, J. See T. Rivoal Sabok, M. and T. Tsankov On the complexity of topological con-

jugacy of Toeplitz subshifts

583

Sahoo, S. K. See J. Chen Sánchez-González, L. See A. Procházka Shalit, O. M. See J. E. M^{c}Carthy Shelah, S. See S. Garti Shelah, S. See M. Malliaris

Sisto, A. Tracking rates of random walks  1

Söze, K. Real zeroes of random polynomials, I. Flip-invariance, Turán’s lemma, and the Newton-Hadamard polygon

817

Söze, K. Real zeroes of random polynomials, II. Descartes’ rule of signs and anti-concentration on the symmetric group

837

Speigelhofer, L. See C. Müller Suomala, V. See T. Ojala Toyoda, M. See K. Aoyama Troitsky, V. G. See N. Gao Tsankov, T. See M. Sabok Umezu, K. See H. R. Quoirin Vázquez-Lorenzo, R. See E. Calviño-Louzao Wang, X. See J. Chen

Wiest, B. On the genericity of loxodromic actions  559

Wilson, B. See D. Bate Wu, M. See T. Ojala Xanthos, F. See N. Gao

Yaskin, V. and N. Zhang Non-central sections of convex bodies  763

Zhang, N. See V. Yaskin